# Actor-Critic Model Predictive Control: Differentiable Optimization meets Reinforcement Learning

### v6

@@ III-G Neural Cost Map
The cost function for the model predictive control architecture presented in Section III-F is learnt as a neural network, depicted in Fig. 1 as Cost Map. Several adaptations to the system are needed in order to properly interface the neural network architecture with the optimization problem. First, we construct the matrix $Q(s_{k})$ and vector $p(s_{k})$ as follows:
@@
The purpose of the diagonalization of the $Q$ matrix is to reduce the dimensionality of the learnable parameter space. Therefore, the dimensionality of the output dimension of the Cost Map is $2T(n_{state}+n_{input})$. In order to ensure the positive semi-definiteness of the $Q$ matrix and the positive definiteness of the R matrix, a lower bound on the value of these coefficients needs to be set. To this end, the last layer of the neural cost map has been chosen to be a sigmoid which allows for upper and lower bounds on the output value. This lower and upper limits are chosen equal for $Q$ and $p$, of 0.1 and 100000.0, respectively. Therefore, the final neural cost map consists of two hidden layers of width 512 with ReLUs in between and a sigmoid non-linearity at the end. The critic network consists also of two hidden layers of width 512 and ReLUs. The ouput of the critic network is a scalar.
@@ V Results
We perform multiple experiments to better understand different properties of our proposed approach. Namely, we study (i) AC-MPC’s performance on multiple drone racing tasks and its sample efficiency in comparison to AC-MLP policies, (ii) its robustness to disturbances, (iii) its robustness to changes in the dynamics, (iv) the benefits of the model-predictive value expansion, (v) the interpretability properties of AC-MPC, (vi) its sensitivity to exploration hyperparameters, (vii) its real-world deployment. Each aspect corresponds to a different research question we attempt to answer in the following.
V-A Does our method improve asymptotic performance and sample efficiency?
@@
In Fig. 2 and in the supplementary video, one can observe the evolution of the value function over time. Given the sparse nature of the reward terms (see Section IV-A3), one can observe that when a gate is successfully passed, the region of high rewards quickly shifts to guide the drone towards the next gate. This can be interpreted as a form of discrete mode switching enabled by the neural network cost map. Such mode-switching behavior is a challenging feat to accomplish using traditional MPC pipelines. The intuition behind this is that the critic is able to learn long-term predictions, while the model-predictive controller focuses on the short-term ones, effectively incorporating two time scales. Additionally, we train our approach on a challenging track known as the SplitS track. Designed by a professional drone racing pilot, this track is distinguished by its highly demanding SplitS maneuver, where the drone must navigate through two gates placed directly above one another successively. This track was used in previous research as a benchmark, namely [14, 22, 9, 10, 11].
@@
In Fig. 3 we show the collective thrust commands of both approaches, AC-MLP vs AC-MPC. In this figure, one can notice how the AC-MPC can achieve saturation in a consistent way, while the AC-MLP approach also saturates but in a more irregular fashion. This indicates that the MPC module within the AC-MPC framework, which includes explicit knowledge of both platform dynamics and its constraints, effectively and consistently utilizes the system’s saturation limits. In contrast, the AC-MLP commands are generated directly by a neural network, resulting in behavior that is comparatively less consistent and less aligned to the platform’s true limits. Additionally, in Fig. 4, we can see the reward evolution for AC-MLP when compared with AC-MPC. This comparison has been conducted after optimizing the initial standard deviation hyperparameter for both approaches, and choosing the best result for each. Generally – and as we study in section V-F – AC-MPC performs better for lower exploration parameter, whereas AC-MLP needs higher exploration. Fig. 4 shows that the asymptotic training performance and sample efficiency are improved when choosing AC-MPC, since it is able to leverage the prior information encoded in the dynamics within the differentiable MPC. This is particularly evident in the reward evolution observed for the Vertical and SplitS tracks. For simpler tasks, the inductive bias introduced in AC-MPC is not strongly advantageous to performance. For instance, in the Horizontal track, which is relatively simple, the training performance of AC-MLP is on par with AC-MPC, since the maneuvers that the policy needs to discover to perform the task do not require an elaborate exploration behavior.
Figure 3: Evolution of collective thrust command after training, for both AC-MLP and AC-MPC in the SplitS track. One can notice how the AC-MPC commands drive the system into a consistent saturation, while the MLP commands are smoothed out in saturation.
@@
Figure 6: Success rate against variations of mass, inertia and body rate limit. Policies are trained in the SplitS track with a nominal value for these parameters. This nominal value is depicted in red. Then, these policies are deployed in 64 paralell environments, each environment having slightly different initial position. In the deployment stage, mass, inertia and body rate limit are varied. The success rate is computed as the proportion of trajectories that don’t have crashes. Because AC-MPC has the dynamics embedded in the network architecture, it shows to be more robust to these changes than its AC-MLP counterpart.
@@ V-E Can we interpret the internal workings of AC-MPC?
Figure 9: Empirical probing of the correlation between the Hessian and the gradient of the learned critic $V(x)$ and the learned differentiable MPC cost terms $Q_{N}$ and $p_{N}$. (A): We train a linear classifier to infer $Q_{N}$ and $p_{N}$ from the Hessian of $V(x)$ and show the prediction from different 2D projections of $Q_{N}$ and $p_{N}$. We show how the groundtruth can be predicted with a high accuracy by only using a small linear classifier, which indicates that the data learned by the critic and $Q_{N}$ and $p_{N}$ are highly correlated. (B): Top: 3D plot of $\tau_{N}^{T}Q_{N}\tau_{N}+\tau_{N}^{T}p_{N}$ for the state dimensions of $v_{x}$ and $v_{y}$ after training. Bottom: 3D plot of $V(x)$ for the same dimensions. Apart from the change in sign and in scale, the similarity between both shapes indicate a high correlation between the value function and the learned cost parameters.
@@
In order to empirically show the relationship between the Hessian of the value function and what is learned in $Q_{N}$ and $p_{N}$, we perform a similarity study using linear probes. Particularly, we first train an AC-MPC agent to perform the stabilization at hover task, introduced in Section V-D. We choose this task here because, since the reward function is quadratic, it results in a optimization landscape that is easier to interpret. Since we need the derivative of the value function with respect to the full state, the critic needs to be trained with full state information. After training, we collect a dataset of uniformly distributed $5242882^{19}=$ 524288 datapoints, where every datapoint corresponds to a sample of $Q_{N}$, $p_{N}$, $H_{V}$ and $\Delta_{V}$, as per Eq. (V-E) and (V-E). Then, with this dataset, we train a single linear layer (wihtout non-linearity) to predict the elements of the learned $Q_{N}$ and $p_{N}$ matrices from the elements of the $H_{V}$ and $p_{V}$, respectively. In the hypothesized case where the information contained in the learned cost is related to the information encoded by the gradient and Hessian of the value function, we expect the regression to have a high accuracy.
@@
In Fig. 9A, we show the true and predicted values of $Q_{N}$ and $p_{N}$. In order to show that the data fully correlates in all dimensions, the data is presented in a an lower triangular matrix for $Q$ and in a upper triangular matrix for $p$. Every row and column of this matrix is composed of the elements of the state space $x$, therefore showing the correlation from different 1-dimensional cuts – for the diagonal elements – and 2-dimensional cuts – for the rest. One can therefore see a high correlation between the predicted and the true values, therefore confirming a strong relationship between the value function and the learned cost parameters. An extra experiment was conducted where we have tried to infer the same Q and p values, but from a random dataset, as a validation that there is no overfitting behaviour happening. Indeed, in this case the linear classifier was not able to predict the true labels at all.
V-F How does our approach handle exploration?
@@
Since our approach incorporates a model of the dynamics within the optimization problem, it has access to prior information before the training phase, and hence leverages an inductive bias in its architecture. This raises questions about the extent of exploration needed to discover an optimal policy and how the exploration parameter can be minimized to achieve optimal performance. To explore this, we study the effect of the initial policy standard deviation hyperparameter on the performance of both AC-MLP and AC-MPC. This hyperparameter influences the exploration behavior of the RL algorithms. In Fig. 10, we present reward plots for the training on the Vertical track under different values of this hyperparameter. The results show that AC-MLP requires a higher degree of exploration to identify the optimal solution. This indicates that AC-MPC effectively utilizes the prior dynamics knowledge, resulting in improved performance with reduced exploration. More interestingly, Fig. 10 shows that AC-MPC is less sensitive to the choice of this hyperparameter and yields high performance despite very low exploration values. On the other hand, AC-MLP fails to reach high-reward regions when trained with low values and generally underperforms in comparison to the proposed method.
@@
V-G How does our method perform in real-world environments?
@@
Additionally, we perform a lap time comparison in the SplitS track. We train both AC-MPC and AC-MLP in the task of drone racing and execute different runs in two different simulation fidelities: nominal dynamics and realistic simulator. This realistic simulator, called NeuroBEM, and based in blade element momentum theory, comes from [122]. In Fig. 11 we show a top view of the deployment of both AC-MLP and AC-MPC in both of the aforementioned simulators. The policies are tested for 64 different initial positions, contained in a cube of $0.5$ m of side length. Additionally, Fig. 12 illustrates trajectories flown in two different tracks (Circle track and SplitS track) in simulation and the real world. Not only can our policy transfer to the real world without any fine-tuning, it can do so while at the limit of handling, being able to achieve speeds of up to $21$ $m/s$. Finally, in Table II, we show the results in terms of lap time and success rate of this comparison, including both simulation modalities, and the real-world results.
@@
As observed in the lap time results, both AC-MPC and AC-MLP exhibit similar performance in terms of lap time, indicating that they are largely on par in this aspect. This similarity is expected, given that their asymptotic performance and final reward values are quite close, as illustrated in Figure 4. However, it is important to note that our approach not only achieves on par performance but also demonstrates additional robustness to out-of-distribution scenarios (Figure 5) and changes in the system dynamics (Figure 6). This indicates that our method maintains its effectiveness under a wider range of conditions and uncertainties, while still matching the performance of the neural network-based architecture.
@@ V-H How do training and inference times compare to standard actor-critic architectures?
In Table II, we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths.
@@ Acknowledgments
We would like to thank Brandon Amos, for sharing his insights regarding the differentiable MPC code. We would also like to thank Jiaxu Xing, Leonard Bauersfeld and Alex Barden for the insightful discussions and support with the hardware setup.

### v7

@@ III-G Neural Cost Map
The cost function for the model predictive control architecture presented in Section III-F is learned as a neural network, depicted in Fig. 1 as Cost Map. Several adaptations to the system are needed in order to properly interface the neural network architecture with the optimization problem. First, we construct the matrix $Q(s_{k})$ and vector $p(s_{k})$ as follows:
@@
The purpose of the diagonalization of the $Q$ matrix is to reduce the dimensionality of the learnable parameter space. Therefore, the dimensionality of the output dimension of the Cost Map is $2T(n_{state}+n_{input})$. In order to ensure the positive semi-definiteness of the $Q$ matrix and the positive definiteness of the R matrix, a lower bound on the value of these coefficients needs to be set. To this end, the last layer of the neural cost map has been chosen to be a sigmoid which allows for upper and lower bounds on the output value. These lower and upper limits are chosen equal for $Q$ and $p$, of 0.1 and 100000.0, respectively. Therefore, the final neural cost map consists of two hidden layers of width 512 with ReLUs in between and a sigmoid non-linearity at the end. The critic network consists also of two hidden layers of width 512 and ReLUs. The ouput of the critic network is a scalar.
@@ V Results
We perform multiple experiments to better understand different properties of our proposed approach. Namely, we study (i) AC-MPC’s performance on multiple drone racing tasks and its sample efficiency in comparison to AC-MLP policies, (ii) its robustness to external disturbances, (iii) its robustness to changes in the dynamics, (iv) the benefits of the Model-Predictive Value expansion, (v) the interpretability properties of AC-MPC, (vi) its sensitivity to exploration hyperparameters, (vii) its real-world deployment. Each aspect corresponds to a different research question that we address in each of the following sections.
V-A Does our method improve performance and sample efficiency?
@@
In Fig. 2 and in the supplementary video, one can observe the evolution of the value function over time. Given the sparse nature of the reward terms (see Section IV-A3), one can observe that when a gate is successfully passed, the region of high rewards quickly shifts to guide the drone towards the next gate. This can be interpreted as a form of discrete mode switching enabled by the neural network cost map. Such mode-switching behavior is a challenging feat to accomplish using traditional MPC pipelines. The intuition behind this is that the critic is able to learn long-term predictions, while the model-predictive controller focuses on the short-term ones, effectively incorporating two time scales.
Additionally, we train our approach on a challenging track known as the SplitS track. Designed by a professional drone racing pilot, this track is distinguished by its highly demanding SplitS maneuver, where the drone must navigate through two gates placed directly above one another successively. This track was used in previous research as a benchmark, namely [14, 21, 9, 10, 11].
@@
In Fig. 3 we show the collective thrust commands of both approaches, AC-MLP vs AC-MPC. In this figure, one can notice how the AC-MPC can achieve saturation in a consistent way, while the AC-MLP approach also saturates but in a more irregular fashion. This indicates that the MPC module within the AC-MPC framework, which includes explicit knowledge of both platform dynamics and its constraints, effectively and consistently utilizes the system’s saturation limits. In contrast, the AC-MLP commands are generated directly by a neural network, resulting in behavior that is comparatively less consistent and less aligned to the platform’s true limits.
In Fig. 4, we can see the reward evolution for AC-MLP when compared with AC-MPC. This comparison was conducted after optimizing the initial standard deviation hyperparameter for both approaches and choosing the best result for each. Generally—as we also study in section V-F—AC-MPC performs better for lower exploration parameters, whereas AC-MLP needs higher exploration parameters. Fig. 4 shows that the asymptotic training performance and sample efficiency are improved when choosing AC-MPC, since it is able to leverage the prior information encoded in the dynamics within the differentiable MPC. This is particularly evident in the reward evolution observed for the Vertical and SplitS tracks. For simpler tasks, the prior knowledge introduced in AC-MPC is not strongly advantageous to performance. For instance, in the Horizontal track, which is relatively simple, the training performance of AC-MLP is on par with AC-MPC, since the maneuvers that the policy needs to discover to perform the task do not require an elaborate exploration behavior.
In conclusion, the results presented in this section show that our method improves both performance and sample efficiency. By comparing AC-MPC against AC-MLP across various track complexities, and after hyperparameter tuning, we have demonstrated that AC-MPC can achieve superior performance and sample efficiency, especially in challenging tasks like the Vertical and SplitS tracks (see Fig. 4). This improvement is attributed to AC-MPC’s ability to effectively leverage prior knowledge of the system dynamics through the embedded MPC module (see Fig. 3). These findings suggest that incorporating model-based control elements into reinforcement learning frameworks offers a promising direction for future research.
Figure 3: Evolution of the collective thrust command over time for both AC-MLP and AC-MPC during trials on the SplitS track. The figure demonstrates AC-MPC’s ability to effectively utilize control input saturation due to its access to the system dynamics and control constraints. In contrast, AC-MLP exhibits less consistent saturation behavior.
@@
Figure 6: Success rate against variations of mass, inertia, and body rate limit. Policies are trained in the SplitS track with a nominal value for these parameters. This nominal value is depicted in red. Then, these policies are deployed in 64 parallel environments, each environment having slightly different initial position. In the deployment stage, mass, inertia, and body rate limit are varied. The success rate is computed as the proportion of trajectories that don’t have crashes. Because AC-MPC has the dynamics embedded in the network architecture, it shows to be more robust to these changes than its AC-MLP counterpart.
@@ V-E Can we interpret the internal workings of AC-MPC?
Figure 9: Empirical probing of the correlation between the Hessian and the gradient of the learned critic $V(x)$ and the learned differentiable MPC cost terms $Q_{N}$ and $p_{N}$. (A): We train a linear classifier to infer $Q_{N}$ and $p_{N}$ from the Hessian of $V(x)$ and show the prediction from different 2D projections of $Q_{N}$ and $p_{N}$. We show how the ground truth can be predicted with high accuracy by only using a small linear classifier, which indicates that the data learned by the critic and $Q_{N}$ and $p_{N}$ are highly correlated. (B): Top: 3D plot of $\tau_{N}^{T}Q_{N}\tau_{N}+\tau_{N}^{T}p_{N}$ for the state dimensions of $v_{x}$ and $v_{y}$ after training. Bottom: 3D plot of $V(x)$ for the same dimensions. Apart from the change in sign and in scale, the similarity between both shapes indicates a high correlation between the value function and the learned cost parameters.
@@
In order to empirically show the relationship between the Hessian of the value function and what is learned in $Q_{N}$ and $p_{N}$, we perform a similarity study using linear probes. Particularly, we first train an AC-MPC agent to perform the stabilization at hover task, introduced in Section V-D. We choose this task here because, since the reward function is quadratic, it results in an optimization landscape that is easier to interpret. Since we need the derivative of the value function with respect to the full state, the critic needs to be trained with full state information. After training, we collect a dataset of uniformly distributed $5242882^{19}=$ 524288 datapoints, where every datapoint corresponds to a sample of $Q_{N}$, $p_{N}$, $H_{V}$ and $\Delta_{V}$, as per Eq. (V-E) and (V-E). Then, with this dataset, we train a single linear layer (without non-linearity) to predict the elements of the learned $Q_{N}$ and $p_{N}$ matrices from the elements of the $H_{V}$ and $p_{V}$, respectively. In the hypothesized case where the information contained in the learned cost is related to the information encoded by the gradient and Hessian of the value function, we expect the regression to have a high accuracy.
@@
In Fig. 9A, we show the true and predicted values of $Q_{N}$ and $p_{N}$. In order to show that the data fully correlates in all dimensions, the data is presented in a lower triangular matrix for $Q$ and in an upper triangular matrix for $p$. Every row and column of this matrix is composed of the elements of the state space $x$, therefore showing the correlation from different 1-dimensional cuts – for the diagonal elements – and 2-dimensional cuts – for the rest. One can therefore see a high correlation between the predicted and the true values, therefore confirming a strong relationship between the value function and the learned cost parameters. An extra experiment was conducted where we have tried to infer the same Q and p values, but from a random dataset, as a validation that there is no overfitting behavior happening. Indeed, in this case the linear classifier was not able to predict the true labels at all.
In conclusion, this section directly investigates and reveals a relationship within the internal workings of AC-MPC. By demonstrating a strong correlation between the Hessian and gradient of the learned value function and the learned cost parameters $Q_{N}$ and $p_{N}$, we have shown that the critic effectively learns information related to the curvature of the cost landscape used by the MPC module. This provides valuable insight into how the critic informs the MPC controller and suggests that the learned value function encodes information about the system’s local dynamics and cost structure.
V-F How sensitive is our approach to exploration hyperparameters?
@@
Since our approach incorporates a model of the dynamics within the optimization problem, it has access to prior information before the training phase, and hence leverages prior knowledge of the dynamics in its architecture. This raises questions about the extent of exploration needed to discover an optimal policy and how the exploration parameter can be minimized to achieve optimal performance. To explore this, we study the effect of the initial policy standard deviation hyperparameter on the performance of both AC-MLP and AC-MPC. This hyperparameter influences the exploration behavior of the RL algorithms. In Fig. 10, we present reward plots for the training on the Vertical track under different values of this hyperparameter. The results show that AC-MLP requires a higher degree of exploration to identify the optimal solution. This indicates that AC-MPC effectively utilizes the prior dynamics knowledge, resulting in improved performance with reduced exploration. More interestingly, Fig. 10 shows that AC-MPC is less sensitive to the choice of this hyperparameter and yields high performance despite very low exploration values. On the other hand, AC-MLP fails to reach high-reward regions when trained with low values and generally underperforms in comparison to the proposed method.
@@
V-G How does our method perform when deployed in the real-world?
@@
Additionally, we perform a lap time comparison in the SplitS track. We train both AC-MPC and AC-MLP in the task of drone racing and execute different runs in two different simulation fidelities: nominal dynamics and realistic simulator. This realistic simulator, called NeuroBEM, and based in blade element momentum theory, comes from [120]. In Fig. 11 we show a top view of the deployment of both AC-MLP and AC-MPC in both of the aforementioned simulators. The policies are tested for 64 different initial positions, contained in a cube of $0.5$ m of side length. Additionally, Fig. 12 illustrates trajectories flown in two different tracks (Circle track and SplitS track) in simulation and the real world. Not only can our policy transfer to the real world without any fine-tuning, but it can do so while at the limit of handling, achieving speeds up to 21 m/s.
In Table II, we show the results of this comparison in terms of lap time and success rate, including both simulation modalities and real-world results.
@@
As observed in the lap time results, both AC-MPC and AC-MLP exhibit similar performance in terms of lap time, indicating that they are largely on par in this aspect. This similarity is expected, given that their asymptotic performance and final reward values are quite close, as illustrated in Figure 4. However, it is important to note that our approach not only achieves on-par performance but also demonstrates additional robustness to out-of-distribution scenarios (see Fig. 5, in Horizontal and Vertical tracks), and changes in the system dynamics (see Fig. 6, in SplitS track). This indicates that our method maintains its effectiveness under a wider range of conditions and uncertainties, while still matching the performance of the neural network-based architecture.
@@ V-H How do training and inference times compare to standard actor-critic architectures?
In Table III, we show the training times (SplitS track) and the forward pass times for AC-MLP and for the proposed AC-MPC for different horizon lengths. As can be seen, while AC-MPC offers performance advantages, it incurs a higher computational cost. Training time is increased by a factor of approximately 30 (for a horizon length of N=2) due to the batched optimization required for each forward and backward pass. This computational overhead also affects online performance, resulting in a slower forward pass. Although the current Python implementation [^1] https://locuslab.github.io/mpc.pytorch/ of the differentiable MPC algorithm [26] achieves a sufficient deployment frequency of 50Hz for our system, optimized C/C++ implementations would significantly reduce training time and enable higher frequency deployments. Note, however, that this computational overhead did not affect closed-loop performance, as the 13.5ms execution time remained well within our 20ms control time step budget necessary to deploy the system successfully in the real world.
@@ Acknowledgments
We would like to thank Brandon Amos, for sharing his insights regarding the differentiable MPC code, and Sebastien Gros, for taking the time to proofread our manuscript. We would also like to thank Jiaxu Xing, Leonard Bauersfeld and Alex Barden for the insightful discussions and support with the hardware setup.